OpenStudy (jamesreed30):

Find the total area for a prism that has sides of 6 4 and eight

OpenStudy (danjs):



OpenStudy (danjs):

The total Surface Area is the sum of all the areas of the sides. calculate each side area and add them all up

OpenStudy (jamesreed30):

So just add 8,6,4 and getvmy answer

OpenStudy (jamesreed30):

@DanJS

OpenStudy (danjs):

You add together the areas of each face. you have 2 that are 6 by 4 you have 2 that are 6 by 8 you have 2 that are 4 by 8

OpenStudy (jamesreed30):

So multiply 6 by 4 twice and etc.?

OpenStudy (jamesreed30):

@DanJS

OpenStudy (danjs):

yeah, 2 sides area 6*4=24 2 sides area 8*6 = 48 2 sides area 8*4 = 32 Total Surface area = 2*(6*4) + 2*(6*8) + 2*(8*4)

OpenStudy (jamesreed30):

I got 104

OpenStudy (jamesreed30):

And also my AND looks like this, (__ +__ squre root symbol)

OpenStudy (jamesreed30):

@DanJS

OpenStudy (danjs):

Total Surface area = 2*(6*4) + 2*(6*8) + 2*(8*4) = 208

OpenStudy (jamesreed30):

Oh I see so how would I do it with the way I showed you my problem is set up

OpenStudy (jamesreed30):

@DanJS